# High-Q cavity coupled to a high permittivity dielectric resonator for sensing applications

### v2

@@ Abstract
The use of high quality factor resonators is of undoubted interest for high precision measurements and for applications in quantum technologies. Novel types of microwave sensors can be realized by coupling a first resonator acting as a stable frequency reference with a second resonator that is sensitive to a particular physical quantity. Here we report on a coupled cavity configuration in which a high $Q$-factor elliptical TESLA-shaped superconducting cavity is coupled with a high permittivity ($\varepsilon_{\mathrm{r}}$) SrTiO3 puck, whose resonant frequency varies as a function of temperature due to the temperature dependence of the permittivity that reaches values higher than 30000 below 1 K. Extensive electromagnetic simulations are used to test different coupling configurations, showing great versatility to tune the coupling in the weak or strong regime, depending on the puck’s position within the cavity. Moreover, for the coupled system, they allow one to investigate the dependence of the zero transmission frequency value on changes of the permittivity $\varepsilon_{\mathrm{r}}$, obtaining a maximum value of 2.8 MHz per unit change of $\varepsilon_{\mathrm{r}}$, for $\varepsilon_{\mathrm{r}}\approx 230$. Finally, we discuss a specific application of the coupled system as a bolometer in different temperature ranges.
pacs:
Valid PACS appear here
preprint: APS/123-QED
@@ I Introduction
Resonant cavities, with their ability to sustain high electromagnetic fields, have emerged as powerful tools for ultra-sensitive measurements. These structures, also known as resonators, operate by confining electromagnetic waves within a bounded region, allowing for resonance at specific frequencies. The high-quality factors ($Q$-factors) associated with resonant cavities provide high sensitivity, and thus potential high measurement precision and accuracy for the measurements of different physical quantities, making them indispensable in various scientific and engineering applications.
The use of resonant cavities for sensitive measurements has a rich history, particularly in the fields of fundamental physics, metrology, and quantum mechanics. In the field of particle physics, high-$Q$ microwave cavities have been instrumental in the search for axions, hypothetical particles proposed as dark matter candidates [1, 2]. Similarly, resonant cavities are used in experiments testing the fundamental constants of nature, providing critical tests of physical theories beyond the Standard Model [3, 4, 5]. In quantum information science, superconducting microwave cavities are employed to couple qubits and facilitate quantum state readout with high fidelity [6, 7]. Furthermore, hybrid quantum systems combining resonant cavities with other quantum technologies, such as optomechanical or spin systems such as nitrogen-vacancy centers in diamond, have shown promise for enhanced sensitivity in detecting weak forces and fields. These hybrid systems leverage the strong coupling between cavity modes and quantum states to achieve unprecedented measurement sensitivity [8]. The inherent sensitivity of resonant cavities also extends to applications in precision spectroscopy and time-keeping [9, 10, 11].
@@
Figure 1: Characterization of the STO resonator at room temperature. (a) Photograph showing the STO puck and Cu cavity used in the experiments. (b) Sketch of the model used for simulations (COMSOL Multiphysics). The top cap is not shown. (c,d) Simulated distribution of the root-mean-square electric and magnetic field for the TE01δ mode ($\varepsilon_{r}=318$). (d,e) Plots of reflection ($S_{11}$) and transmission spectra ($S_{21}$) measured at room temperature (incident power 0 dBm). The amplitude is shown in blue, the phase in green. (f) Comparison between simulated and experimental $S_{21}$ spectra. The peak at 1.22 GHz is reproduced by the the simulated $\mathrm{TE_{01\delta}}$ mode, while the dip displayed by the simulation at $\approx 1\,\mathrm{GHz}$ is below the background transmission of the cavity ($-100\,\mathrm{dB}$). The peak at 1.8 GHz is probably related to the hybrid $\mathrm{HEM_{12\delta}}$ mode, although in this case the simulation shows a mismatch of $\approx 200\,\mathrm{MHz}$.
@@
The coupling of two, or more, resonators has been exploited for several of the above-mentioned applications [12, 13, 14, 15, 16, 17]. If a high $Q$ factor resonator provides a stable reference frequency, the coupling with a second resonator can be used as an additional resource and to make the system sensitive to a specific physical quantity. For example, superconducting cavities were coupled in different ways showing that the photon lifetime of the local field can exceed that of the bare cavities [18]. Coupled sapphire resonators showed increased $Q$-factor and tunability due to the supermode effect, which was studied for axion detection with high sensitivity [19]. Moreover, systems composed of coupled dielectric resonators were investigated for the bolometric detection of particles and X-ray photons [12, 20, 21, 22]. There, a sapphire resonator acting as a stable reference frequency is coupled to a second resonator having a strongly temperature-dependent permittivity (e.g. CaTiO3 or SrTiO3). The absorption of energy causes a change in the absorber’s permittivity, in turn producing a measurable change in the resonant frequency of the coupled system. Additionally, coupled resonator systems typically show a transmission zero, or notch, at the frequency for which the injected power is completely reflected back. The notch frequency and depth are sensitive probes of the frequency and coupling of the resonant modes [23, 24, 17].
@@
Dielectric resonators have attracted long time interest for several applications in sensing, metrology and detectors due to their well defined resonant modes and high $Q$-factors. High permittivity materials such as SrTiO3, KTaO3, LaAlO3 and TiO2 have been used in different situations, including room-temperature masers [25], dark matter searches [26] quantum sensing [27] and spin manipulation [28]. In particular, strontium titanate (SrTiO3) has been the subject of intense study because of its peculiar physics [29], characterized in particular by quantum paraelectricity [30, 31, 32] and superconductivity [33] at low temperature, multiferroicity [34] and emergent conductivity or magnetism at the interface with other oxides, i.e. in LaAlO3/SrTiO3 heterostructures [29]. The properties of SrTiO3 at microwave frequencies were also investigated [35, 36, 37], recently reporting also the anomalous behavior of complex permittivity below 1 K [38, 39].
@@
Here we focus on the investigation of a particular system of coupled resonators, in which a strontium titanate (hereafter STO) puck is placed in a three-dimensional elliptical cavity. The latter specifically corresponds to a superconducting TESLA cavity at 1.3 GHz, whose geometry has been developed in the context of particle accelerators to obtain $Q$-factors as high as $10^{9}$ [40]. We first characterize the STO resonator alone and test the evolution of permittivity and losses at temperatures down to 150 mK. We then exploit finite-element electromagnetic simulations to study the coupling between elliptical cavity and dielectric resonator, discussing the tuning of frequency and $Q$-factor of the coupled modes as a function of geometric and material parameters, such as the position of the STO puck within the cavity and the permittivity of the STO puck. These results demonstrate great tunability for the proposed coupled system and allow the definition of conditions to control the coupling between high-$Q$ cavities and high permittivity dielectric resonators, opening a path for the exploitation of these systems in applications. In particular, we discuss the exploitation of coupled cavity-STO resonator system for the realization of bolometers working in different temperature ranges.
@@ II Characterization of the STO resonator
The dielectric resonator was realized with a STO crystal produced by SurfaceNet GmbH that was shaped into a cylinder having radius $a$=8.17 mm and height $d$=7.26 mm, with a central hole of radius 2 mm. For the characterization of the dielectric resonator we used a hollow copper cavity (inner radius 60 mm, height 22 mm), whose lowest mode at 3.8 GHz stands well above the fundamental mode of the dielectric resonator in order to avoid cross couplings between their modes. The STO puck was installed on top of two quartz spacers and coupled to the coaxial lines by loop antennae positioned at a distance of 10-20 mm from the outer surface of the cylinder (Fig. 1(a)). Measurements were carried out with the whole set-up installed either at room temperature, in liquid nitrogen or in a dilution fridge.
@@
Reflection and transmission spectra acquired at room temperature show the fundamental mode of the STO resonator at frequency $f_{0}=1.22$ GHz (Fig. 1(d,e)). Considering the geometric factors of the resonator and the STO permittivity $\epsilon_{r}\approx 318$, we obtain that $f_{0}$ corresponds well with the frequency of the $\mathrm{TE_{01\delta}}$ mode derived by the semi-analytical formula [41]
$f_{STO}\approx\frac{\alpha}{\sqrt{\varepsilon_{\mathrm{r}}}}$ (1)
being $\alpha\mathrm{[GHz]}=34/(a\mathrm{[mm]})\left(a/d+3.45\right)$, while $a$ and $d$ are the radius and height of the puck, respectively. The loaded quality factor extracted from the transmission spectrum (Fig. 1(e)) is $Q_{L}=8000$. Since the loops are positioned far apart from the dielectric puck, the small coupling between feedlines and resonator turns into insertion loss of $30\,\mathrm{dB}$. Thus, the internal quality factor can be approximated as $Q_{0}\approx Q_{L}\approx 1/\tan{\delta}$. The loss tangent of STO results $\tan{\delta}=1.2\times 10^{4}$. Finite-element electromagnetic simulations confirm the distribution expected for the $\mathrm{TE_{01\delta}}$ mode (Fig. 1(c,d)), whose frequency results about 1.2 GHz in close agreement with experimental data (Fig. 1(f)).
@@
The temperature evolution between 4 and 300 K of cylindrical STO resonators is reported in the literature [35, 37, 36]. In this temperature range, the maximum $Q$-factor value of 20000 is reached around 100 K, just above the structural temperature transition of STO. The nonlinear change in the permittivity of STO encompasses two orders of magnitude to obtain $\varepsilon_{r}\approx 3\times 10^{4}$ and $\tan{\delta}\approx 1\times 10^{-4}$ at 5 K [35].
We have tested the STO resonator at temperatures below 1 K (Fig. 2). The STO resonator was slowly cooled down to 150 mK in about 6 hours. The temperature was measured using a thermometer installed into the copper cavity. At the lowest temperature, the $\mathrm{TE_{01\delta}}$ mode frequency resulted $f_{0}=116.9$ MHz with $Q\approx 1\times 10^{4}$; these values suggest $\varepsilon_{r}\approx 3\times 10^{4}$ and $\tan{\delta}\approx 1\times 10^{-4}$. As a function of the decreasing temperature, the frequency of the $\mathrm{TE_{01\delta}}$ mode decreases until 0.4 K, while for lower temperatures we observe a slight upturn. In this temperature ($T$) range, the frequency of the resonator can be reproduced using a polynomial curve, $f_{STO}=K_{0}+K_{1}T+K_{2}T^{2}$. The fit of the experimental points (Fig. 2(b)) gives $K_{0}=116.88\leavevmode\nobreak\ \mathrm{MHz}$, $K_{1}=-0.008\leavevmode\nobreak\ \mathrm{MHz/K}$ and $K_{2}=0.012\leavevmode\nobreak\ \mathrm{MHz/K^{2}}$.
@@
III Coupled STO and elliptical cavity
This section examines the coupling of the STO puck with a resonating cavity to form a coupled resonator system. To facilitate this investigation, we selected an elliptical TESLA-shaped cavity[40] operating at 1.3 GHz. This frequency was selected because it closely matches the operational frequency of the available STO puck sample at room temperature and because similar elliptical cavity prototypes were available for measurement at the European Organization for Nuclear Research (CERN). However, it is important to emphasize that the choice of cavity and frequency is not limited to the elliptical cavity at 1.3 GHz; other cavity types and frequencies may be used based on specific project requirements. The primary aim of this study is to assess the feasibility of coupling between the STO puck and a superconducting RF cavity.
Electromagnetic simulations.
The TESLA-shaped elliptical cavity and its fundamental mode (FM), which is a transverse magnetic (TM) mode, are illustrated in Fig. 3. The cavity’s equator radius was adjusted to achieve a frequency of 1.3 GHz. For a copper cavity, the quality factor of this mode at room temperature is $2.89\times 10^{4}$. By assuming $\varepsilon_{r}=318$, the first mode of the STO puck has a frequency of 1.10 GHz and a quality factor of $1.01\times 10^{4}$. The field plot of this mode is also shown in Fig. 3. The eigenmode and frequency domain simulations in this section were performed using CST Studio Suite.
@@ III Coupled STO and elliptical cavity
To investigate the coupling between the two modes, the STO puck is placed inside the elliptical cavity. Two essential parameters for effective coupling are the orientation of the STO puck within the elliptical cavity and its radial offset from the center. The puck’s position in the elliptical cavity for effective coupling is shown in Fig. 4. Coupling is examined using both eigenmode analysis and frequency domain analysis, with two identical antennas, each having $Q_{\mathrm{ext}}=8.6\times 10^{7}$. In the eigenmode analysis, no antennas are considered, and the only losses accounted for are the dielectric losses of STO and surface losses on the elliptical cavity if non-perfect electric conductor (PEC) materials are used.
Figure 5: Parameter sweep of the dielectric constant ($\varepsilon_{r}$) of STO and its effect on the frequency and quality factor of the first two modes of the coupled elliptical cavity and STO puck at different radial offset ($r_{\mathrm{offset}}$) values. Mode 1 is the mode with the lower frequency. (a) Frequency of the eigenmodes, (b) quality factor of the eigenmodes, and (c) E-field distributions of the two modes at $\varepsilon_{\mathrm{r}}=230$ and $r_{\mathrm{offset}}=50$ mm.
@@
The results of the eigenmode analysis are presented in Fig. 5. The investigation examines how varying $\varepsilon_{r}$ of the STO, which can be a function of temperature, affects the resonant frequency and $Q$-factor of the first two modes in the coupled system. The STO-dominant mode is characterized by having most of its energy localized in the STO, while the elliptical cavity-dominant mode retains most of its energy within the elliptical cavity, similar to the uncoupled system. Below $\varepsilon_{r}=230$, Mode 2 is the STO-dominant mode and Mode 1 is the elliptical cavity-dominant mode. Above $\varepsilon_{r}=230$, the order of the modes is reversed. When $r_{\mathrm{offset}}=$0 mm, the two modes do not interact as $\varepsilon_{r}$ changes. In this scenario, the elliptical cavity’s frequency (Fig. 5(a)) and $Q$-factor (Fig. 5(b)) remain nearly constant at 1.3 GHz and $Q_{\mathrm{ext}}=4.2\times 10^{7}$, respectively, for a PEC cavity. In this PEC situation, the $Q$-factor value is influenced by the losses in the dielectric region occupied by the STO puck. Conversely, the frequency of the STO-dominant mode decreases with increasing $\varepsilon_{r}$, while its $Q$-factor remains nearly constant at $Q_{\mathrm{ext}}=1.0\times 10^{4}$. In this scenario, the frequencies of the elliptical cavity-dominant mode at 1.3 GHz and the STO-dominant mode are unaffected by each other and intersect at $\varepsilon_{r}=230$, with no interaction between the modes as $\varepsilon_{r}$ changes (Fig. 5(a)). The permittivity value of 230 corresponds to a temperature above room temperature, even though the system is ideally intended for operation at lower temperatures where the STO’s resonance frequency exhibits the greatest sensitivity to temperature variations (around 25 K). However, due to the fixed dimensions of the STO puck and the elliptical cavity’s frequency, we adjust the STO permittivity to shift its resonance frequency near 1.3 GHz to investigate the mode coupling between the two resonators. Alternative dielectrics having low temperature permittivity values closer to those considered in the simulations were also considered. In particular, a rutile puck having similar dimensions shows, at 0.8 K, a mode at 2.6 GHz with $Q$-factor of $5.6\times 10^{5}$ [42]. However the variation of the permittivity of rutile as a function of temperature is less marked [43], thus coupled resonator systems based on rutile resonators are expected to be less sensitive as bolometers.
When the STO puck is radially offset, mode splitting occurs at $\varepsilon_{r}=230$, and the interaction between the two modes intensifies as the radial offset increases from zero. The level repulsion, defined as the distance between the two dispersion curves, grows with larger radial offsets, indicating stronger coupling between the two modes (Fig. 5(a)). This coupling causes the $Q$-factor of the STO-dominant mode (low-$Q$ mode) to increase, while the $Q$-factor of the elliptical cavity-dominant mode (high-$Q$ mode) decreases, and both converge at $\varepsilon_{r}=230$ (Fig. 5(b)). For a PEC elliptical cavity and an STO puck with $r_{\mathrm{offset}}$=50 mm and $\varepsilon_{r}=230$, the two modes exhibit equal quality factors of approximately $Q=2\times 10^{4}$. One mode has a frequency of 1279 MHz, while the other is at 1315 MHz. The field plots for these two modes are presented in Fig. 5(c). In addition to changes in $\varepsilon_{r}$, the eigenmode frequency sensitivities and their quality factors also depend on the radial offset. As shown in the Fig. 6, at $\varepsilon_{r}$ = 230, where the coupling is at its maximum, even a slight radial offset of the puck results in a sharp reduction in the quality factor of the elliptical cavity-dominant mode. This is followed by an increase in the mode’s frequency sensitivity to changes in $\varepsilon_{r}$.

### v3

@@ Abstract
The use of high-quality factor resonators is of undoubted interest for high-precision measurements and for applications in quantum technologies. Novel types of microwave sensors can be realized by coupling a first resonator acting as a stable frequency reference with a second resonator that is sensitive to a particular physical quantity. Here we report on a coupled cavity configuration in which a high $Q$ factor elliptical TESLA-shaped superconducting cavity is coupled with a high permittivity ($\varepsilon_{\mathrm{r}}$) SrTiO3 puck, whose resonant frequency varies as a function of temperature due to the temperature dependence of the permittivity that reaches values higher than 30000 below 1 K. Extensive electromagnetic simulations are used to test different coupling configurations, showing great versatility in tuning the coupling in the weak or strong regime, depending on the puck’s position within the cavity. Moreover, for the coupled system, they allow investigation of the dependence of the zero transmission frequency (notch frequency) on changes in permittivity $\varepsilon_{\mathrm{r}}$, obtaining a maximum value of 2.8 MHz per unit change of $\varepsilon_{\mathrm{r}}$, for $\varepsilon_{\mathrm{r}}\approx 230$. Finally, we discuss the use of the coupled system as a sensor that operates in different temperature ranges.
@@ I Introduction
Resonant cavities, with their ability to sustain high electromagnetic fields, have emerged as powerful tools for ultrasensitive measurements. These structures, also known as resonators, operate by confining electromagnetic waves within a bounded region, allowing resonance at specific frequencies. The high quality factors ($Q$ factors) associated with resonant cavities provide high sensitivity and therefore potential high measurement precision and accuracy for measurements of different physical quantities, making them indispensable in various scientific and engineering applications.
@@
Figure 1: Characterization of the STO resonator at room temperature. (a) Photograph showing the STO puck and Cu cavity used in the experiments. (b) Sketch of the model used for simulations (COMSOL Multiphysics). The top cap is not shown. (c,d) Simulated distribution of the root-mean-square electric and magnetic field for the TE01δ mode ($\varepsilon_{r}=318$). (d,e) Plots of reflection ($S_{11}$) and transmission spectra ($S_{21}$) measured at room temperature (incident power 0 dBm). The amplitude is shown in blue, the phase in green. (f) Comparison between simulated and experimental $S_{21}$ spectra. The peak at 1.22 GHz is reproduced by the the simulated $\mathrm{TE_{01\delta}}$ mode, while the dip displayed by the simulation at $\approx 1\,\mathrm{GHz}$ is below the background transmission due to the noise floor of the measurement ($100\,\mathrm{dB}$). The peak at 1.8 GHz is probably related to the hybrid $\mathrm{HEM_{12\delta}}$ mode, although in this case the simulation shows a mismatch of $\approx 200\,\mathrm{MHz}$.
The use of resonant cavities for sensitive measurements has a rich history, particularly in the fields of fundamental physics, metrology, and quantum mechanics. For example, early work by Pound and Rebka (1959) demonstrated the potential of resonant cavities to detect extremely small frequency shifts, paving the way for their use in precision measurement [1]. In the field of particle physics, high-$Q$ microwave cavities have been instrumental in the search for axions, hypothetical particles proposed as dark matter candidates [2, 3]. Similarly, resonant cavities are used in experiments testing the fundamental constants of nature, providing critical tests of physical theories beyond the Standard Model [4, 5, 6]. In quantum information science, superconducting microwave cavities are employed to couple qubits and facilitate the quantum state readout with high fidelity [7, 8]. Furthermore, hybrid quantum systems combining resonant cavities with other quantum technologies, such as optomechanical or spin systems such as nitrogen-vacancy centers in diamond, have shown promise for enhanced sensitivity in detecting weak forces and fields. These hybrid systems leverage the strong coupling between cavity modes and quantum states to achieve unprecedented measurement sensitivity [9]. The inherent sensitivity of resonant cavities also extends to applications in precision spectroscopy and time-keeping [10, 11, 12].
@@
The coupling of two or more resonators has been exploited for several of the applications mentioned above [13, 14, 15, 16, 17, 18, 19, 20, 21]. If a high $Q$ factor resonator provides a stable reference frequency, the coupling with a second resonator can be used as an additional resource and to make the system sensitive to a specific physical quantity. For example, superconducting cavities were coupled in different ways, showing that the photon lifetime of the local field can exceed that of the bare cavities [22]. Coupled sapphire resonators showed an increase in the $Q$ factor and tunability due to the supermode effect, which was studied for axion detection with high sensitivity [23]. Moreover, systems composed of coupled dielectric resonators were investigated for the bolometric detection of particles and X-ray photons [13, 24, 25, 26]. There, a sapphire resonator that acts as a stable reference frequency is coupled to a second resonator that has a strongly temperature-dependent permittivity (e.g. CaTiO3 or SrTiO3). Absorption of energy causes a change in the permittivity of the absorber, which in turn produces a measurable change in the resonant frequency of the coupled system. Additionally, coupled resonator systems typically show a transmission zero, or notch, at the frequency for which the injected power is completely reflected back. The frequency and depth of the notch are sensitive probes of the frequency and coupling of the resonant modes [27, 28, 21].
@@
Dielectric resonators have attracted long-time interest for several applications in sensing, metrology, and detectors due to their well-defined resonant modes and high $Q$ factors. High permittivity materials such as SrTiO3, KTaO3, LaAlO3 and TiO2 have been used in different situations, including room temperature masers [29], dark matter searches [30] quantum sensing [31] and spin manipulation [32]. In particular, strontium titanate (SrTiO3) has been the subject of intense study due to its peculiar physics [33], characterized in particular by quantum paraelectricity [34, 35, 36] and superconductivity [37] at low temperature, multiferroicity [38] and emergent conductivity or magnetism at the interface with other oxides, i.e. in LaAlO3/SrTiO3 heterostructures [33]. The properties of SrTiO3 microwave resonators were also investigated [39, 40, 41, 42], showing permittivity values reaching 30000 at 10 K and recently reporting an anomalous behavior below 1 K [43, 44].
@@
Here we report the study of a particular system of coupled resonators, in which a strontium titanate (hereafter STO) puck, whose resonant frequency is strongly dependent on temperature or other external parameters, is placed in a three-dimensional high-$Q$ cavity. In particular, we consider as a prototype an elliptical TESLA-shaped cavity with fundamental mode at 1.3 GHz, whose geometry has been developed in the context of particle accelerators and optimized for manufacture with superconducting materials to obtain $Q$ factors as high as $10^{9}$ [45, 46], with potential applications also in dark matter searches [47] and quantum computing [48]. We first characterize the STO resonator alone and test the evolution of permittivity and losses at temperatures down to 160 mK. We then exploited finite element electromagnetic simulations to study the coupling between the elliptical cavity and the dielectric resonator, discussing the tuning of frequency and $Q$ factor of the coupled modes as a function of geometric and material parameters, such as the position of the STO puck within the cavity and the permittivity of the STO puck. These results demonstrate great tunability for the proposed coupled system and allow the definition of conditions to control the coupling between high-$Q$ cavities and high-permittivity dielectric resonators, opening a path for the exploitation of these systems in applications. In particular, we discuss the use of the coupled cavity-STO resonator system in a bolometer operating at different temperatures.
@@ II Characterization of the STO resonator
The dielectric resonator was realized with a STO crystal produced by SurfaceNet GmbH that was shaped into a cylinder having radius aa=8.17 mm and height dd=7.26 mm, with a central hole of radius 2 mm. For the characterization of the dielectric resonator we used a hollow copper cavity (inner radius 60 mm, height 22 mm), whose lowest mode at 3.8 GHz stands well above the fundamental mode of the dielectric resonator in order to avoid cross-couplings between their modes. The STO puck was installed on top of two quartz spacers and coupled to the coaxial lines by loop antennae positioned at a distance of 10-20 mm from the outer surface of the cylinder (Fig. 1(a)). Measurements were carried out with the whole set-up installed at room temperature, in liquid nitrogen, or in a dilution fridge.
@@
Reflection and transmission spectra acquired at room temperature show the fundamental mode of the STO resonator at frequency $1.22$ GHz (Fig. 1(d,e)). Considering the geometric factors of the resonator and the STO permittivity $\epsilon_{r}\approx 318$, we find that this number corresponds well to the frequency of the $\mathrm{TE_{01\delta}}$ mode derived by the semi-analytical formula [49]
$f_{\mathrm{STO}}\approx\frac{\alpha}{\sqrt{\varepsilon_{\mathrm{r}}}}\$(1)
being $\alpha\mathrm{[GHz]}=34/(a\mathrm{[mm]})\left(a/d+3.45\right)$, while $a$ and $d$ are the radius and height of the puck, respectively. The loaded quality factor extracted from the transmission spectrum (Fig. 1(e)) is $Q_{L}=8000$. Since the loops are located far apart from the dielectric puck, the small coupling between the feedlines and the resonator results in an insertion loss of $30\,\mathrm{dB}$. Thus, the internal quality factor can be approximated as $Q_{0}\approx Q_{L}\approx 1/\tan{\delta}$. The loss tangent of the STO results $\tan{\delta}=1.2\times 10^{4}$. Finite element electromagnetic simulations confirm the expected distribution for the $\mathrm{TE_{01\delta}}$ mode (Fig. 1(c,d)), whose frequency results about 1.2 GHz in close agreement with the experimental data (Fig. 1(f)).
@@
The temperature evolution between 4 and 300 K of cylindrical STO resonators is reported in the literature [39, 40, 42, 41]. In this temperature range, the maximum value of the $Q$ factor of 20000 is reached around 100 K, just above the structural temperature transition of the STO. The nonlinear change in the permittivity of STO encompasses two orders of magnitude to obtain $\varepsilon_{r}\approx 3\times 10^{4}$ and $\tan{\delta}\approx 1\times 10^{-4}$ at 5 K [40].
We tested the STO resonator at temperatures below 1 K (Fig. 2). The STO resonator was slowly cooled to 160 mK in about 6 hours. The temperature was measured using a thermometer installed in the copper cavity. At the lowest temperature, the frequency of the $\mathrm{TE_{01\delta}}$ mode resulted in $116.9$ MHz with $Q\approx 1\times 10^{4}$; these values suggest $\varepsilon_{r}\approx 3\times 10^{4}$ and $\tan{\delta}\approx 1\times 10^{-4}$. As a function of the decreasing temperature, the frequency of the $\mathrm{TE_{01\delta}}$ mode decreases until 0.4 K, while for lower temperatures we observe a slight upturn. In this temperature ($T$) range, the frequency of the resonator can be reproduced using a polynomial curve, $f_{STO}=K_{0}+K_{1}T+K_{2}T^{2}$. The fit of the experimental points (Fig. 2(b)) gives $K_{0}=116.88\leavevmode\nobreak\ \mathrm{MHz}$, $K_{1}=-0.008\leavevmode\nobreak\ \mathrm{MHz/K}$ and $K_{2}=0.012\leavevmode\nobreak\ \mathrm{MHz/K^{2}}$.
@@
Figure 5: Parameter sweep of the dielectric constant ($\varepsilon_{r}$) of STO and its effect on the frequency and quality factor of the first two modes of the coupled elliptical cavity and STO puck at different radial offset ($r_{\mathrm{offset}}$) values. Mode 1 is the mode with the lower frequency. (a) Frequency of the eigenmodes, (b) quality factor of the eigenmodes, and (c) E-field distributions of the two modes at $\varepsilon_{\mathrm{r}}=230$ and $r_{\mathrm{offset}}=50$ $m$.
@@
III Coupled STO and elliptical cavity
This section examines the coupling of the STO puck with a resonating cavity to form a coupled resonator system. To facilitate this investigation, we selected an elliptical TESLA-shaped cavity[45] operating at frequency $f_{cav}=$1.3 GHz. This frequency was selected because it closely matches the operational frequency of the available STO puck sample at room temperature and because similar elliptical cavity prototypes were available for measurement at the European Organization for Nuclear Research (CERN). However, it is important to emphasize that the choice of cavity and frequency is not limited to the elliptical cavity at 1.3 GHz; other cavity types and frequencies can be used based on specific project requirements. The primary aim of this study is to assess the feasibility of coupling between the STO puck and a superconducting RF cavity.
Electromagnetic simulations.
The TESLA-shaped elliptical cavity and its fundamental mode (FM), which is a transverse magnetic (TM) mode, are illustrated in Fig. 3. The radius of the equator of the cavity was adjusted to achieve a frequency of 1.3 GHz. For a copper cavity, the quality factor of this mode at room temperature is $2.89\times 10^{4}$. Assuming $\varepsilon_{r}=318$, the first mode of the STO puck has a frequency of 1.10 GHz and a quality factor of $1.01\times 10^{4}$. The field plot of this mode is also shown in Fig. 3. The eigenmode and frequency domain simulations in this section were performed using the CST Studio Suite.
To investigate the coupling between the two modes, the STO puck is placed inside the elliptical cavity. Two essential parameters for effective coupling are the orientation of the STO puck within the elliptical cavity and its radial offset from the center. The position of the puck in the elliptical cavity for effective coupling is shown in Fig. 4. The coupling is examined using both the eigenmode analysis and the frequency domain analysis, with two identical antennas, each having $Q_{\mathrm{ext}}=8.6\times 10^{7}$. In the eigenmode analysis, no antennas are considered, and the only losses accounted for are the dielectric losses of STO and surface losses on the elliptical cavity if non-perfect electric conductor (PEC) materials are used.
The results of the eigenmode analysis are presented in Fig. 5. The investigation examines how the variation $\varepsilon_{r}$ of the STO, which can be a function of temperature, affects the resonant frequency and the $Q$ factor of the first two modes in the coupled system. The STO-dominant mode is characterized by having most of its energy localized in the STO, while the elliptical cavity-dominant mode retains most of its energy within the elliptical cavity, similar to the uncoupled system. Below $\varepsilon_{r}=230$, Mode 2 is the STO-dominant mode and Mode 1 is the elliptical cavity-dominant mode. Above $\varepsilon_{r}=230$, the order of the modes is reversed. When $r_{\mathrm{offset}}=$0 mm, the two modes do not interact as $\varepsilon_{r}$ changes. In this scenario, the elliptical cavity’s frequency (Fig. 5(a)) and $Q$ factor (Fig. 5(b)) remain nearly constant at 1.3 GHz and $Q_{\mathrm{ext}}=4.2\times 10^{7}$, respectively, for a PEC cavity. In this PEC situation, the value of the $Q$ factor is influenced by the losses in the dielectric region occupied by the STO puck. In contrast, the frequency of the STO-dominant mode decreases with increasing $\varepsilon_{r}$, while its $Q$ factor remains nearly constant at $Q_{\mathrm{ext}}=1.0\times 10^{4}$. In this scenario, the frequencies of the elliptical cavity-dominant mode at 1.3 GHz and the STO-dominant mode are not affected by each other and intersect in $\varepsilon_{r}=230$, without interaction between the modes as $\varepsilon_{r}$ changes (Fig. 5(a)). The permittivity value of 230 corresponds to a temperature above room temperature, although the system is ideally intended for operation at lower temperatures where the STO’s resonance frequency exhibits the greatest sensitivity to temperature variations (around 25 K). However, due to the fixed dimensions of the STO puck and the elliptical cavity’s frequency, we adjusted the STO permittivity to shift its resonance frequency near 1.3 GHz to investigate the mode coupling between the two resonators. Alternative dielectrics with low-temperature permittivity values closer to those considered in the simulations were also considered. In particular, a rutile puck having similar dimensions shows, at 0.8 K, a mode at 2.6 GHz with $Q$ factor of $5.6\times 10^{5}$ [50]. However, the variation in the permittivity of rutile as a function of temperature is less marked [51], thus coupled resonator systems based on rutile resonators are expected to be less sensitive as bolometers.
When the STO puck is radially offset, mode splitting occurs at $\varepsilon_{r}=230$, and the interaction between the two modes intensifies as the radial offset increases from zero. Level repulsion, defined as the distance between the two dispersion curves, increases with larger radial offsets, indicating stronger coupling between the two modes (Fig. 5(a)). This coupling causes the $Q$ factor of the STO-dominant mode (low $Q$ mode) to increase, while the $Q$ factor of the elliptical cavity-dominant mode (high $Q$ mode) decreases, and both converge at $\varepsilon_{r}=230$ (Fig. 5(b)). For a PEC elliptical cavity and an STO puck with $r_{\mathrm{offset}}$=50 mm and $\varepsilon_{r}=230$, the two modes exhibit equal quality factors of approximately $Q=2\times 10^{4}$. One mode has a frequency of 1279 MHz, while the other is at 1315 MHz. The field plots for these two modes are presented in Fig. 5(c). In addition to changes in $\varepsilon_{r}$, the frequency sensitivities of the eigenmodes and their quality factors also depend on the radial offset. As shown in Fig. 6, at $\varepsilon_{r}$ = 230, where the coupling is at its maximum, even a slight radial offset of the puck results in a sharp reduction in the quality factor of the elliptical cavity-dominant mode. This is followed by an increase in the frequency sensitivity of the mode to changes in $\varepsilon_{r}$.